# DXAI: Explaining Classification by Image Decomposition

### v1

@@ Appendix C More results
In the main paper we presented qualitative results of our algorithm for different datasets with selected comparisons to other methods. Here we provide more detail. We show for each image the full comparison to 6 established XAI methods. We also provide more examples in each dataset, so the reader can judge better the quality and stability of our algorithm. We show the results for the case $C=D$ where the classifier is the discriminator (where classification accuracy is best) and for $C=ResNet18$, cases which were not shown in the main paper. In Fig. 17 we show comparison between Female and Male classes. There are many subtle cues for such distinction. One can observe that our method provides dense informative explanations which highlight well the distinctions. In Fig. 18 the peppers are shown, which differ mainly by color, as our proposed method shows clearly. Heatmap visualization cannot show well global color explanations. In Fig. 19 we show that the cars are well isolated in high resolution by our method. This case is less dense and can be handled well also by some other methods (such as LRP, Guided grad-cam or Gradient SHAP). Our method appears to visualize the explanation very clearly. In Fig. 20 and Fig. 21 we show additional datasets of tomatoes and apples classes, which did not appear in the main paper. Here again color changes are a main feature of distinction, which is shown clearly by our method. In Fig. 22 additional examples of the BraTS dataset are provided, showing our ability to isolate the lesions well. In Fig. 23 we observe the dense distinction images of animal classes, where different fur textures are significant features. These are very hard to visualize by heatmaps. The above images were using the discriminator as the classifier. In Fig. 24, Fig. 25, Fig. 26 and Fig. 27 we show results for ResNet18 as the classifier. We can expect that since both classifiers are relatively advanced, with high precision, they consider similar features. We can observe that our method produce qualitatively quite similar results and is stable. For the peppers case, Fig. 27, we observe a stronger difference, where the agnostic part is more of a mix of orange and red colors. In any case, the color distinction is clear and is much better visualized than in standard XAI methods. In Figs 28, 29, 30, 31 the classification explanations for the Simple classifier are shown. One can observe that the distinction map is less accurate, more blurry, containing some spurious background features. This can be expected for a lower quality classifier.

### v2

@@ Appendix 0.G More results
In the main paper we presented qualitative results of our algorithm for different datasets with selected comparisons to other methods. Here we provide more details. We show for each image the full comparison to 6 established XAI methods. We also provide more examples in each dataset, so the reader can judge better the quality and stability of our algorithm. We show the results for the case $C=D$ where the classifier is the discriminator (where classification accuracy is best) and for $C=ResNet18$, cases which were not shown in the main paper. In Fig. 15 we show comparison between Female and Male classes. There are many subtle cues for such distinction. One can observe that our method provides dense informative explanations which highlight well the distinctions. In Fig. 16 the peppers are shown, which differ mainly by color, as our proposed method shows clearly. Heatmap visualization cannot show well global color explanations. In Fig. 17 we show that the cars are well isolated in high resolution by our method. This case is less dense and can be handled well also by some other methods (such as LRP, Guided grad-cam or Gradient SHAP). Our method appears to visualize the explanation very clearly. In Fig. 18 and Fig. 19 we show additional datasets of tomatoes and apples classes, which did not appear in the main paper. Here again color changes are a main feature of distinction, which is shown clearly by our method. In Fig. 20 additional examples of the BraTS dataset are provided, showing our ability to isolate the lesions well. In Fig. 21 we observe the dense distinction images of animal classes, where different fur textures are significant features. These are very hard to visualize by heatmaps. The above images were using the discriminator as the classifier. In Fig. 22, Fig. 23, Fig. 24 and Fig. 25 we show results for ResNet18 as the classifier. We can expect that since both classifiers are relatively advanced, with high precision, they consider similar features. We can observe that our method produce qualitatively quite similar results and is stable. For the peppers case, Fig. 25, we observe a stronger difference, where the agnostic part is more of a mix of orange and red colors. In any case, the color distinction is clear and is much better visualized than in standard XAI methods. In Figs 26, 27, 28, 29 the classification explanations for the Simple classifier are shown. One can observe that the distinction map is less accurate, more blurry, containing some spurious background features. This can be expected for a lower quality classifier.